# On the (In)Approximability of the Monitoring Edge Geodetic Set Problem

### v1

@@ 5.2.2 Algorithm description
We start by arguing that $M_{r}$ achieves the claimed approximation ratios for $H_{r}$. When $\beta>0$, this follows directly from Lemma 5.16 and from the fact that $G_{r}$ has $n_{r}=c(V)$ vertices. Regarding the case $\beta=0$, if $u$ is the parent of $v$, then $n_{v}\leq(1-\alpha)n_{u}$. Then, the tree $T$ has height at most $\left\lceil\log_{\frac{1}{1-\alpha}}n_{r}\right\rceil=\left\lceil\frac{\log c(V)}{\log\frac{1}{1-\alpha}}\right\rceil$, and Lemma 5.18 implies that the achieved approximation is at most $\left(1+\left\lceil\frac{\log c(V)}{\log\frac{1}{1-\alpha}}\right\rceil\right)\cdot\sqrt{2\beta_{0}\ln n}$.
To obtain the claimed approximations for the input instance of GMEG-Set, let $Z$ be the set of vertices having cost 00 in such instance, and consider $M=M_{r}\cup Z$. Since $c(M)=c(M_{r})$, $c(V)=c(V(G^{\prime}))$, and $n\geq n_{r}$, the claim on the approximation trivially holds, and we only need to argue that $M$ is indeed a GMEG-Set.
By Lemma 5.11, $M_{r}$ monitors, in $G$, all edges of $E^{\prime}$ that belong to $V(G^{\prime})\cup\delta V(G^{\prime})$. Hence, if $V(G^{\prime})\cup\delta V(G^{\prime})=V(G)$, we are done. Otherwise, let $A=V(G)\setminus(V(G^{\prime})\cup\delta V(G^{\prime}))$, and let $H_{A}$ be the projection of $A$. Since $A\subseteq Z$, the set $Z$ is a GMEG-Set for $H_{A}$, and by Lemma 5.11, $Z$ monitors, in $G$, all edges of $E^{\prime}$ that belong to $V(A)\cup\delta V(A)$. Hence, $A\cup M_{r}\subseteq Z\cup M_{r}$ monitors all edges in $E^{\prime}$, i.e., it is a GMEG-Set.

### v3

@@ 5.2.2 Algorithm description
We start by arguing that $M_{r}$ achieves the claimed approximation ratios for $H_{r}$. When $\beta>0$, this follows directly from Lemma˜5.16 and from the fact that $G_{r}$ has $n_{r}=c(V)$ vertices. Regarding the case $\beta=0$, if $u$ is the parent of $v$, then $n_{v}\leq(1-\alpha)n_{u}$. Then, the tree $T$ has height at most $\left\lceil\log_{\frac{1}{1-\alpha}}n_{r}\right\rceil=\left\lceil\frac{\log c(V)}{\log\frac{1}{1-\alpha}}\right\rceil$, and Lemma˜5.18 implies that the achieved approximation is at most $\left(1+\left\lceil\frac{\log c(V)}{\log\frac{1}{1-\alpha}}\right\rceil\right)\cdot\sqrt{2\beta_{0}\ln n}$.
To obtain the claimed approximations for the input instance of GMEG-Set, let $Z$ be the set of vertices having cost 0 in such instance, and consider $M=M_{r}\cup Z$. Since $c(M)=c(M_{r})$, $c(V)=c(V(G^{\prime}))$, and $n\geq n_{r}$, the claim on the approximation trivially holds, and we only need to argue that $M$ is indeed a GMEG-Set.
By Lemma˜5.11, $M_{r}$ monitors, in $G$, all edges of $E^{\prime}$ that belong to $V(G^{\prime})\cup\delta V(G^{\prime})$. Hence, if $V(G^{\prime})\cup\delta V(G^{\prime})=V(G)$, we are done. Otherwise, let $A=V(G)\setminus(V(G^{\prime})\cup\delta V(G^{\prime}))$, and let $H_{A}$ be the projection of $A$. Since $A\subseteq Z$, the set $Z$ is a GMEG-Set for $H_{A}$, and by Lemma˜5.11, $Z$ monitors, in $G$, all edges of $E^{\prime}$ that belong to $V(A)\cup\delta V(A)$. Hence, $A\cup M_{r}\subseteq Z\cup M_{r}$ monitors all edges in $E^{\prime}$, i.e., it is a GMEG-Set.